OpenStudy (anonymous):

College basketball is played on a court that is 94 feet long and 50 feet wide. The high school team has to do laps around the court's edge. What is the distance around the court? 268 feet 288 feet 318 feet 248 feet

OpenStudy (lukaryse):

Distance around the court=Perimeter Perimeter is solved by P=2w+2l, where w is width and l is length We know the width and the length, so we put it in the formula P=2(50)+2(94) P=100+188 P=288